payton:

Maya had $27. She spent all the money buying three burgers for x dollars each and two sandwiches for y dollars each. If Maya had bought two burgers and one sandwich, she would have been left with $11. The following system of equations models this scenario: 3x + 2y = 27 2x + y = 16 Use the system of equations to solve for x and y. (6, 5) (5, 6) (3, 2) (2, 3)

ramen:

You can graph this one if you want.

ramen:

But I'll let imagine finish

payton:

no it would be 17

payton:

no that would be 18

payton:

ohh i read it like the last one my bad

Shadow:

You can eyeball this but that's kind of cheating. \[2x + y = 16\] \[y = -2x + 16\] \[3x + 2y = 27\] \[3x + 2(-2x + 16) = 27\] \[3x - 4x + 32 = 27\] \[-x + 32 = 27\] \[-x = -5 \rightarrow x = 5\] Then just input x = 5 into either equation to solve for y

payton:

yes it would be 16

justus:

3x+2y=27 2x+y=16 now multiply by -2 -2(2x+y)=-2*16 which is 4x-2y=-32 after we can add the equations together 3x+2y=27 -4x-2y=-32 3x-4x=27-32 -x=-5 x=5 now we substitute the x 2(5)+y=16 10+y=16 y=6 therefore the solution set is (5,6)

ramen:

Yup plug the y into the y of the other equation.

carmelle:

do u need help with anything else?